Assuming a​ 1-year, money market account investment at 3.09 percent​ (APY), a 2.01​% inflation​ rate, a 15 percent marginal tax​ bracket, and a constant $60,000 ​balance, calculate the​ after-tax rate of​ return, the real return and the total monetary return. What are the implications of this result for cash management​ decisions?
Answer #1

real rate of return= 1+ nominal rate of return/ 1+ inflation rate

= 1+3.09/1+2.01= 1.3588

tax bracket is 15%, hence the after tax return= (1- tax rate) interest rate= 1-0.15*1.3588= 1.155

total monetary return on $60,000= $60,000*1.155%= $693 is the monetary return.

The present investment resulting in positive cash inflows to the investor. While we are taking decisions for investing money in fixed income securities, we must consider inflation rates, interest rates and also liquidity to the securities. Perhaps this investment alternative is providing positive returns from the investment and suggest to go with investment.